Please help me solve: (2n + 1)/7 >= (n + 4)/5

Answer by Maricel Balos(8)

(2n + 1)/7 >= (n + 4)/5
35[(2n+1)/7 >= [(n+4)/5]
5(2n+1) >= 7(n+4)
10n+5 >= 7n+28
10n-7n >= 28-5
3n >= 23
3n/3 >= 23/3
n >= 7 2/3